OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

geerky42 (geerky42):

Use slope-point formula: \(\large y-y_1 = m(x-x_1)\)

OpenStudy (anonymous):

y+4=3/5(x-2)

OpenStudy (anonymous):

my problem is the fraction, what to do.

geerky42 (geerky42):

just do same as if slope is integer? you just need equation